Hi, not sure if this question is in the correct area but please can some one answer it?
Solve the following equation:
3t - 12= 4- t
Prove your solution is correct
Can I please have all the working out, it helps me understand how to complete the question.
This is much appreciated

Answer by Cromlix(4381)

3t - 12 = 4 - t
Collect like terms.
Remember to change signs of
unknowns or numbers if you cross
the = sign.
3t + t = 4 + 12
4t = 16
Divide both sides by 4
t = 4.
Hope this helps.
:-)